OpenStudy (aravindg):

lim x tends to 0+ (sin x)^x=??

OpenStudy (aravindg):

\[\lim_{x \rightarrow 0+} ( sinx)^{x}=??\]

OpenStudy (aravindg):

.......

OpenStudy (anonymous):

1

OpenStudy (jamesj):

Well, \[\lim_{x \rightarrow 0+} x^x = 1\] and we know that sin x / x --> 1 as x --> 0, so intuitively we would expect that \[\lim_{x \rightarrow 0+} (\sin x)^x = 1\] Now to prove it.

OpenStudy (aravindg):

??how?

OpenStudy (aravindg):

so

OpenStudy (jamesj):

First take the ln of the expression and we have \[\ln( (\sin x)^x ) = x \ln(\sin x) = { \ln(\sin x) \over 1/x}\] Now numerator and denominator go to zero as x --> 0+. Let's apply l'Hopital's rule; then this last term is equal to \[{ \cos x / \sin x \over -1/x^2} = x \cos x . {x \over \sin x}\]

OpenStudy (jamesj):

Now the limit of each of these three terms respectively is 0, 1 and 1 hence \[\lim_{x \rightarrow 0+} \ln((\sin x)^x) = 0\]

OpenStudy (jamesj):

sorry, dropped the minus sign, but the result is the same.

OpenStudy (jamesj):

Now. If \[\lim_{x \rightarrow 0+} f(x) = A\] then \[lim_{x \rightarrow 0+} e^{f(x)} = e^A

OpenStudy (jamesj):

\[lim_{x \rightarrow 0+} e^{f(x)} = e^A \]

OpenStudy (jamesj):

and so you're done.

OpenStudy (aravindg):

srry i am not allowed to use l hospital rule

OpenStudy (jamesj):

In that case, bound sin x above and below by some functions, f(x) and g(x) such that \[\lim_{x \rightarrow 0+} f(x)^x = \lim_{x \rightarrow 0+} g(x)^x = 1.\] For example, for 0 < x < 1, \[\sin x \leq x \]

OpenStudy (jamesj):

and we know that x^x has limit 1 as x -->0+

OpenStudy (jamesj):

Now you need a lower bound for sin x, g(x) where you can calculate the limit of g(x)^x

OpenStudy (aravindg):

oh cmplicated

OpenStudy (jamesj):

For example g(x) = x/10 will work. I.e., for 0 < x < 1 \[{x \over 10} \leq \sin x \leq x\] hence \[\left({x \over 10}\right)^x \leq (\sin x)^x \leq x^x\]

OpenStudy (jamesj):

and limit of both functions on either side as x --> 0+ is 1. Hence the limit we want is 1.

OpenStudy (jamesj):

what happened to your picture btw?

OpenStudy (aravindg):

?

OpenStudy (jamesj):

your profile pic

OpenStudy (jamesj):

it was good. Anyway, this is how I think you have to calculate your limit.

OpenStudy (aravindg):

isnt the pic good?

OpenStudy (jamesj):

the new one? It's fine.

OpenStudy (jamesj):

But the old one was better if you ask me. But anyway.

OpenStudy (aravindg):

:)

OpenStudy (turingtest):

hey James can you help me see where I messed up? The question if from sonofa_nh, you can scroll down to find it.

OpenStudy (aravindg):

\[\lim_{x \rightarrow 0} e ^{(1/x*\log x)}\]

OpenStudy (aravindg):

=??

OpenStudy (jamesj):

post as new question

OpenStudy (aravindg):

k

OpenStudy (anonymous):

@James: I am not sure, but should we necessarily need to apply L'hospital here? \[ \ln((\sin x)x)=x \ln(\sin x) \] now, applying the limits to zero we directly could write the whole thing as zero,since anything multiplied to zero gives zero Isn't?

OpenStudy (turingtest):

but ln(0) is undefined, so no, right

OpenStudy (jamesj):

No, because ln(sin x) --> -infinity and we don't know for sure that the x is going to cancel it out fast enough

OpenStudy (anonymous):

Assuming you know that: \[ \lim_{x\to0} \frac{\sin{x}}{x} = 1 \quad \text{and} \quad \lim_{x\to0^{+}} x \ln{x} = 0\] You can solve it easier. First note that: \[(\sin{x})^{x} = \exp{(\ln{(\sin{x})^x})} = \exp{(x \ln{(\sin{x})})} \] Then multiply the exponent with (sin x / sin x) resulting in: \[ \exp{(\frac{x}{\sin{x}} \sin{x} \ln{(\sin{x})})} \to \exp{(1\times0)} = e^{0} =1 \quad \text{when} \quad x \to 0^{+} \]